### v1

# Understanding Emotion in Discourse: Recognition Insights and Linguistic Patterns for Generation

Abstract
While Emotion Recognition in Conversation (ERC) has achieved high accuracy, two critical gaps remain: a limited understanding of which architectural choices actually matter, and a lack of linguistic analysis connecting recognition to generation. We address both gaps through a systematic analysis of the IEMOCAP dataset.
For recognition, we conduct a rigorous ablation study with 10-seed evaluation and report three key findings. First, conversational context is paramount, with performance saturating rapidly—90% of the total gain achieved within just the most recent 10–30 preceding turns (depending on the label set). Second, hierarchical sentence representations help at utterance-level, but this benefit disappears once conversational context is provided, suggesting that context subsumes intra-utterance structure. Third, external affective lexicons (SenticNet) provide no gain, indicating that pre-trained encoders already capture necessary emotional semantics. With simple architectures using strictly causal context, we achieve 82.69% (4-way) and 67.07% (6-way) weighted F1, outperforming prior text-only methods including those using bidirectional context.
For linguistic analysis, we analyze 5,286 discourse marker occurrences and find a significant association between emotion and marker positioning ($p<.0001$). Notably, sad utterances exhibit reduced left-periphery marker usage (21.9%) compared to other emotions (28–32%), consistent with theories linking left-periphery markers to active discourse management. This connects to our recognition finding that sadness benefits most from context (+22%p): lacking explicit pragmatic signals, sad utterances require conversational history for disambiguation.
1 Introduction
Emotion recognition in conversation (ERC) is a central challenge for building socially intelligent dialogue systems, mental health support tools, and empathetic AI agents. Unlike sentence-level emotion detection, ERC requires models to account for conversational context, speaker roles, and subtle pragmatic cues. Recent advances with large pretrained encoders have achieved impressive accuracy—state-of-the-art text-only methods reach 81.4% (4-way) and 64.4% (6-way) weighted F1 on IEMOCAP (dutta2024hcam). Yet these empirical gains come at the cost of transparency: we have little understanding of what architectural choices actually drive performance.
Gap 1: What actually matters?
The field has pursued increasingly complex architectures: hierarchical attention mechanisms (majumder2019dialoguernn), knowledge graph integration (zhong2020ket), external lexicon fusion (tu2022context), and elaborate context modeling (ghosal2019dialoguegcn). Yet we lack systematic understanding of which components genuinely contribute. Most prior studies report single-seed results without statistical tests, making it difficult to distinguish genuine improvements from random variation. Do we really need hierarchical sentence representations? Does external affective knowledge help, or do pretrained encoders already capture this information? How much conversational context is sufficient?
Gap 2: Recognition does not inform generation.
High classification accuracy does not translate to understanding how emotions manifest linguistically. Current ERC models provide no guidance for generating emotionally appropriate text. What linguistic patterns distinguish sad from angry utterances? The role of discourse markers in expressing subjectivity and intersubjectivity is well-established in linguistics (schiffrin1987discourse; beeching2014discourse), yet their positional patterns in ERC remain unexplored. Bridging this gap could benefit both recognition (through linguistically-motivated features) and generation (through actionable production guidelines).
1.1 Research Questions
We address these gaps through systematic empirical analysis on IEMOCAP, combining rigorous ablation studies with large-scale discourse analysis. Our investigation is guided by four research questions:
Recognition: What architectural choices matter?
  • RQ1
    Does conversational context improve emotion recognition, and how much is sufficient?
  • RQ2
    Does hierarchical sentence representation help?
  • RQ3
    Does external affective lexicon (SenticNet) help?
Linguistic Analysis: What patterns exist?
  • RQ4
    Are there emotion-specific discourse marker patterns that could inform generation?
1.2 Contributions
  1. 1.
    We conduct the first ERC study with 10-seed evaluation, paired $t$-tests, and Bonferroni correction, establishing a standard for reliable comparison.
  2. 2.
    We isolate the contributions of context, hierarchical structure, and external lexicons, revealing that context is paramount while lexicons provide no benefit.
  3. 3.
    We show that different emotions require different amounts of context—Sad benefits most (+22%p) while Angry benefits least (+8%p)—challenging arousal-based explanations.
  4. 4.
    We conduct the first large-scale analysis of discourse marker positioning in ERC (5,286 occurrences), revealing that Sad utterances show significantly reduced left-periphery usage.
  5. 5.
    We achieve 82.69% (4-way) and 67.07% (6-way) using strictly causal context, outperforming bidirectional methods and enabling real-time deployment.
2 Related Work
2.1 Emotion Recognition in Conversation
Early ERC methods focused on capturing conversational dynamics through recurrent architectures. DialogueRNN (majumder2019dialoguernn) models speaker states across turns, achieving 76.2% on IEMOCAP 4-way classification. COSMIC (ghosal2020cosmic) incorporates commonsense knowledge for context enhancement (77.4%). Recent work has pursued increasingly complex architectures: graph-based models (ghosal2019dialoguegcn), hierarchical attention (ma2022multi), and multimodal transformers (hu2022unimse).
Among text-only methods, HCAM (dutta2024hcam) achieves 81.4% (4-way) and 64.4% (6-way) through hierarchical Bi-GRU with self-attention, while EmoCaps (li2022emocaps) reaches 69.49% on 6-way using emotion capsule networks. Critically, both methods exploit bidirectional context—future utterances unavailable in real-time settings. Our strictly causal approach achieves 82.69% on 4-way, surpassing all text-only methods including bidirectional ones, and 67.07% on 6-way, outperforming six bidirectional methods despite using only past context.
A critical limitation across these methods is their arbitrary context modeling: most fix KK=3 preceding utterances or process entire conversations without considering that different emotions may require different context lengths. While zhang2023variable introduced adaptive instance-level context selection, they still treat all emotions uniformly. We address this through systematic context analysis, exploring whether different emotions exhibit distinct saturation patterns.
2.2 Discourse Markers and Affective Lexicons
Linguistic theory has long recognized discourse markers as signals of speaker stance and emotion (schiffrin1987discourse). fraser1999discourse classified pragmatic markers by function, while beeching2014discourse demonstrated their tendency to cluster at utterance peripheries. The left periphery typically hosts subjective markers expressing speaker attitude: well signals hesitation, oh marks surprise or realization, and actually introduces contrast or correction (traugott2010intersubjectivity).
Despite extensive theoretical study, computational models have largely ignored these positional patterns. Prior ERC models treat all token positions equally, missing potential emotional cues encoded in discourse structure. We address this gap through systematic analysis of discourse markers in IEMOCAP, examining whether positional patterns vary by emotion. To our knowledge, this is the first large-scale corpus study linking discourse marker positions to emotion categories in conversational speech.
Affective lexicons provide complementary emotional knowledge. While early resources like WordNet-Affect (strapparava2004wordnet) offered categorical labels, SenticNet (cambria2024senticnet) maps concepts onto psychological dimensions. Knowledge-enhanced models (zhong2020ket; tu2022senticgat) show improvements, but whether lexicons add unique information beyond contextual encoders remains unclear—a question we address empirically.
3 Methodology
(a) Dialogue length distribution
(b) Sentence distribution
(c) 4-way emotion patterns
(d) 6-way emotion patterns
Figure 1: IEMOCAP dataset characteristics.
3.1 Task and Dataset
We conduct experiments on the IEMOCAP dataset (busso2008iemocap), which contains approximately 12 hours of audiovisual data from 10 actors performing scripted and improvised dyadic conversations across 110 dialogue sessions. The dataset exhibits substantial variability in dialogue length (Figure 1a), with sessions ranging from 24 to 267 utterances (mean: 91.7, median: 67.5). Approximately 70% of dialogues contain fewer than 100 utterances, while the 95th percentile reaches 196 utterances, highlighting the long-tail distribution inherent in conversational data.
Turn-level context length and K-sweep protocol.
We define the context length $K$ as the number of preceding turns included as strictly past-only conversational history for a target utterance. Because there is no established stopping rule for how much prior dialogue is sufficient in ERC, we perform an exhaustive bottom-up sweep $K\in\{0,1,\ldots,K_{\max}\}$ to characterize the performance–context trade-off and identify saturation behavior. We set $K_{\max}$ to the maximum dialogue length (in turns) observed in the corpus distribution used to construct dialogue histories. Importantly, this sweep is used strictly for analysis (saturation characterization) and not for test-set model selection; we do not choose $K$ to maximize test performance.
For a fair comparison with state-of-the-art models (dutta2024hcam), we follow the standard speaker-disjoint split strategy: Sessions 2–4 serve as training data (6,072 utterances from 68 dialogues), Session 1 as validation (1,819 utterances from 20 dialogues), and Session 5 as test set (2,196 utterances from 22 dialogues). This split ensures speaker independence between training and evaluation, with each dialogue containing an average of 89–100 utterances across splits. The sentence-level structure within utterances (Figure 1b) shows that most utterances contain 2–5 sentences, motivating our comparison of flat versus hierarchical encoding strategies.
We evaluate on two emotion taxonomies commonly used in the literature. The 4-way classification task considers angry (1,103 utterances, 19.9%), happy (1,636, 29.6%), sad (1,084, 19.6%), and neutral (1,708, 30.9%) emotions, where the excited category is merged into happy (Figure 1c). The 6-way task extends this to include excited (1,041, 14.1%) and frustrated (1,849, 25.1%) as separate categories (Figure 1d). We use weighted F1-score as our primary metric to account for class imbalance inherent in conversational emotion data. We supplement this with class-wise F1 scores and confusion matrices for detailed error analysis.
3.2 Discourse Marker Analysis
Beyond recognition accuracy, we conduct a linguistic analysis of discourse markers (DMs) to examine emotion-specific pragmatic patterns that may inform future work on emotion-conditioned dialogue generation. DMs are lexical expressions that signal relationships between discourse segments rather than contributing to propositional content (schiffrin1987discourse; fraser1999discourse). Drawing from established taxonomies (schiffrin1987discourse; fraser1999discourse; traugott2010intersubjectivity; beeching2014discourse), we identify 20 markers occurring in IEMOCAP, including turn-management markers (well, oh), connectives (and, but, so), and stance markers (I think, I guess, maybe, you know, I mean). See Appendix A for the full inventory and frequency distribution.
For each marker occurrence, we record its relative position within the utterance (normalized to $[0,1]$), its periphery classification, and the emotion label of the containing utterance. In discourse, the left periphery (LP) and right periphery (RP) serve asymmetric functions (for a comprehensive review, see beeching2014discourse): LP is where speakers claim the floor and manage topic structure, serving textual and subjective functions, while RP is oriented toward the hearer, serving intersubjective and modalising functions. We operationally define LP as position <0.15<0.15, RP as position >0.85>0.85, and medial otherwise. For consistency with our pooling notation, wmean_pos_rev emphasizes utterance-initial (left-peripheral) tokens, whereas wmean_pos emphasizes utterance-final (right-peripheral) tokens; mean pooling treats all positions uniformly.
To test for emotion-specific positional patterns, we employ ANOVA to compare mean positions across emotions, $\chi^{2}$ tests with Cramér’s $V$ to assess association between periphery categories and emotions, mixed-effects models with dialogue as random intercept to control for dialogue-level variation, and post-hoc pairwise comparisons with Bonferroni correction.
[…]
Figure 2: Model architecture. Each turn is encoded independently via Sentence-RoBERTa using either flat (whole utterance) or hierarchical (sentence-level) encoding. SenticNet features are optionally fused (dashed boxes). For classification, we use MLP when K=0 (no context) or unidirectional LSTM when K>0 (with preceding turns as context).
3.3 Model Architecture
To investigate how emotional information is encoded and propagated in dialogue, we develop two encoder variants: a flat encoder and a hierarchical encoder. The flat encoder processes each utterance as a single sequence, while the hierarchical encoder first encodes individual sentences within an utterance, then aggregates them to form the utterance representation. This allows us to examine whether intra-utterance structure provides additional benefit for emotion recognition.
During embedding generation, we include all utterances from each dialogue, even those without target emotion labels (i.e., labels other than happy, sad, angry, neutral for 4-way classification) or unlabeled utterances. This captures long-range contextual dependencies while excluding these utterances from classification training and evaluation. Utterances within each dialogue are indexed to preserve sequential order for turn-level context modeling. Figure 2 illustrates our architecture.
Utterance Encoder.
We compare two encoding strategies: (1) flat encoding, which treats each utterance (turn) as a single sequence and extracts the pooled representation, and (2) hierarchical encoding, which first encodes individual sentences within an utterance, then aggregates them into an utterance-level representation.
We evaluate three pre-trained encoders: BERT-base-uncased (devlin2019bert), RoBERTa-base (liu2019roberta), and Sentence-RoBERTa (NLI-RoBERTa-base-v2) (reimers-2019-sentence-bert). For layer selection, we compare two strategies: (1) avg_last4, which averages the last four transformer layers (layers 9–12), capturing high-level semantic features while avoiding overspecialization of the final layer, and (2) last, which uses only the final layer output.
Classifier.
For utterance-level classification (K=0), the utterance representation is passed through a two-layer MLP with ReLU activation and dropout. When incorporating turn-level context (K>0), we process the sequence of utterance representations through a single-layer unidirectional LSTM and use the final hidden state for classification.
Lexical Feature Integration.
To examine whether external affective knowledge benefits emotion recognition, we integrate SenticNet 7 (cambria2024senticnet), which provides four-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude) for words and phrases. For each utterance, we extract SenticNet features by matching constituent words, aggregate them via mean pooling, and concatenate with the encoder representation. This allows us to test whether pre-trained encoders already capture affective semantics or whether explicit lexical knowledge provides additional benefit.
3.4 Training and Implementation Details
Our classifier consists of a two-layer MLP (for K=0) or a single-layer unidirectional LSTM (for K>0). We use fixed hyperparameters across all experiments: learning rate of 1e-3, hidden dimension of 256, dropout rate of 0.3, and batch size of 64. Training employs Adam optimizer with early stopping based on validation loss (patience of 60 epochs for utterance-level, 20 epochs for turn-level experiments).
All experiments use 10 random seeds {42, 43, …, 51} with results reported as mean $\pm$ standard deviation. Implementation uses PyTorch 1.13 with experiments conducted on NVIDIA A100 GPUs via Saturn Cloud.
4 Experiments and Analysis
4.1 Encoder Selection
We evaluated three pre-trained encoders (Table 1). Sentence-RoBERTa achieves the best performance (65.29%), likely due to its NLI fine-tuning which implicitly captures affective reasoning. All subsequent experiments use Sentence-RoBERTa.
Table 1: Encoder comparison on 4-way classification (K=0, 10 seeds).
Encoder WF1 (%) Std Min Max 95% CI
BERT-base 63.99 0.85 62.62 65.18 [63.38, 64.59]
RoBERTa-base 65.01 0.80 63.90 66.31 [64.44, 65.58]
Sentence-RoBERTa 65.29 1.17 64.10 67.31 [64.45, 66.12]
4.2 Main Results
Table 2 presents our main results comparing flat and hierarchical encoding strategies. A striking pattern emerges: at utterance-level (K=0), hierarchical encoding outperforms flat encoding (+3.17%p for 4-way, $p<.01$; +1.80%p for 6-way, $p<.05$), suggesting that intra-utterance sentence structure aids emotion recognition. However, when conversational context is incorporated, flat encoding achieves the best performance (82.69% for 4-way, 67.07% for 6-way), though the difference from hierarchical encoding is not statistically significant ($p=.427$ for 4-way, $p=.078$ for 6-way). This reversal suggests that turn-level context subsumes the structural information captured by hierarchical encoding. We note that variations in layer selection (last vs. avg_last4) and pooling methods showed no significant differences (14 comparisons, all $p>.08$; see Appendix), justifying our reporting of only the best configurations.
Table 2: Main results on IEMOCAP (10 seeds). $K$ denotes the number of preceding turns (past-only context). Each row shows the best configuration for that encoding strategy. Hierarchical encoding helps at utterance-level (K=0), but flat encoding wins with conversational context.
Task K Encoding Layer Pool Mean $\pm$ Std Min Max 95% CI
4-way 0 FLAT last wmean_pos 65.29 $\pm$ 1.17 64.10 67.31 [64.45, 66.12]
4-way 0 HIER avg_last4 wmean_pos_rev 68.46 $\pm$ 1.09 66.30 70.12 [67.68, 69.24]
4-way 132 FLAT avg_last4 mean 82.69 $\pm$ 0.50 81.73 83.43 [82.33, 83.05]
4-way 132 HIER avg_last4 mean 81.89 $\pm$ 0.41 81.22 82.60 [81.60, 82.18]
6-way 0 FLAT last wmean_pos 52.69 $\pm$ 1.04 50.61 54.51 [51.95, 53.43]
6-way 0 HIER avg_last4 wmean_pos_rev 54.49 $\pm$ 0.84 53.17 55.77 [53.88, 55.09]
6-way 101 FLAT avg_last4 wmean_pos 67.07 $\pm$ 0.69 65.75 68.04 [66.58, 67.57]
6-way 101 HIER avg_last4 mean 66.73 $\pm$ 0.87 65.53 68.68 [66.11, 67.36]
4.3 Emotion-Specific Context Effects
While prior work on variable-length context focuses on how to adaptively select context windows through speaker-aware modules (zhang2023emotion), we take a complementary approach: we systematically investigate what patterns emerge when varying context length and whether different emotions exhibit distinct context requirements.
To analyze emotion-specific effects, we compute per-class F1 scores at each context length from K=0 (utterance only) to K=200 (preceding turns). For each emotion, we report the context length at which F1 attains its maximum within the sweep range, and quantify improvement relative to $K=0$. We also determine the saturation point, defined as the minimum K at which 90% of maximum improvement is achieved. All analyses are conducted across 10 random seeds.
Our first finding is that performance saturates rapidly. As shown in Figure 3, 90% of maximum improvement is achieved with K=30 for 4-way classification and K=10 for 6-way classification. These saturation points correspond to a small fraction of typical dialogue history measured in turns. This suggests that the immediate preceding context carries most of the predictive information for emotion recognition, and sophisticated adaptive selection mechanisms may offer diminishing returns for typical utterances.
Our second finding reveals a dissociation between saturation timing and improvement magnitude. While all emotions reach saturation at similar context lengths (Kruskal-Wallis $H=0.52$, $p=0.91$; mean saturation K ranges from 43 to 61 across emotions), the magnitude of improvement differs dramatically across emotions (one-way ANOVA $F=136.80$, $p<0.0001$). Table 3 presents the full breakdown. For 4-way classification, Sad shows the largest improvement (+22.31 percentage points), followed by Happy (+15.82%p), Neutral (+10.77%p), and Angry (+8.34%p). This ordering is remarkably consistent in 6-way classification, where Sad again benefits most (+21.58%p) and Angry benefits least (+9.43%p), with Frustrated showing similarly low improvement (+8.73%p).
The context length at which each emotion attains its peak performance varies. For 4-way classification, Angry peaks at $K=10$ (10 preceding turns), whereas Sad, Happy, and Neutral peak at K=130K=130140140 (130–140 preceding turns). In 6-way classification, Sad peaks at $K=200$, while Angry peaks at $K=60$. These patterns suggest that certain emotions are recognizable from immediate local cues, while others depend on accumulated conversational dynamics.
We discuss the theoretical implications of these findings, including why arousal alone cannot explain the observed patterns and what this reveals about the linguistic expression of different emotions, in Section 5.
Figure 3: Per-emotion F1 scores across context sizes (4-way classification). Angry saturates quickly at $K^{*}=10$ (10 preceding turns), while Sad, Happy, and Neutral require extended context ($K^{*}=130$140140 preceding turns). Shaded regions indicate $\pm 1$ standard deviation across 10 seeds.
[…]
4-way Classification
[…]
Sad 46.17 130 +22.31
Happy 58.53 140 +15.82
Neutral 43.92 140 +10.77
Angry 35.27 10 +8.34
6-way Classification
[…]
Sad 46.19 200 +21.58
Excited 51.68 90 +10.99
Neutral 43.23 100 +10.69
Happy 31.73 110 +9.68
Angry 41.11 60 +9.43
Frust. 46.66 90 +8.73
4.4 Ablation Studies
We conducted ablation studies on pooling strategy, external lexical knowledge, and layer selection.
For pooling strategy, we compared mean pooling, position-weighted pooling emphasizing utterance-final tokens (wmean_pos), and position-weighted pooling emphasizing utterance-initial tokens (wmean_pos_rev). No method achieved statistically significant superiority (Friedman test, $p>.08$). Position-weighted pooling yielded the best configurations at utterance-level (K=0), while mean pooling performed best with conversational context (K>0).
For external lexical knowledge, we integrated SenticNet, which provides 4-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude). However, this yielded no improvement (4-way: 0.94-0.94%; 6-way: $\sim$0%), suggesting that pre-trained language model embeddings already capture sufficient affective semantics (see Appendix C for complete results across 36 configurations).
For layer selection, averaging the last four transformer layers (avg_last4) versus using only the final layer (last) showed no significant difference (paired $t$-tests, min $p=.244$).
4.5 Comparison with Prior Work
We compare our approach against prior text-only methods on IEMOCAP. A critical distinction among ERC methods is their temporal context access: bidirectional methods utilize both past and future utterances, while past-only (causal) methods access only preceding context. This distinction is crucial for real-time deployment scenarios where future utterances are unavailable.
4-way Classification.
Table 4 shows that our past-only approach achieves the highest performance among all text-only methods, including those using bidirectional context. We outperform HFFN by +1.15%p and HCAM by +1.29%p, despite these methods having access to future context. This result demonstrates that our discourse-aware pooling and emotion-specific context modeling capture patterns that bidirectional architectures miss.
[…]
Method Context WF1 (%)
Ours Past-only 82.69
HFFN (mai2019divide) Bidirectional 81.54
HCAM (dutta2024hcam) Bidirectional 81.4
CHFusion (majumder2018multimodal) Bidirectional 73.6
6-way Classification.
Table 5 presents results on the more challenging 6-way task. Among past-only methods, DAG-ERC achieves 68.03%, while our mean performance across 10 seeds is 67.07%. While we report mean performance across 10 seeds for statistical rigor, our best run achieves 68.04%, marginally surpassing DAG-ERC’s reported 68.03%. More notably, our past-only approach outperforms six bidirectional methods, demonstrating that access to future context does not guarantee superior performance when discourse-aware features effectively capture emotional dynamics from past context alone.
The 2.42%p gap between EmoCaps and our method can be attributed to EmoCaps’ bidirectional context and emotion capsule architecture specifically designed for capturing emotional tendencies. Nevertheless, the competitive performance of our simpler, strictly causal approach suggests that carefully designed discourse features can partially compensate for the absence of future context.
Table 5: Comparison on IEMOCAP 6-way classification (text-only methods, excluding LLM-based approaches). Bold indicates best past-only result.
Method Context WF1 (%)
EmoCaps (li2022emocaps) Bidirectional 69.49
DAG-ERC (shen2021directed) Past-only 68.03
Ours Past-only 67.07
SKAIG (li2021past) Bidirectional 66.96
DialogueCRN (hu2021dialoguecrn) Bidirectional 66.20
CoG-BART (li2022cogbart) Bidirectional 66.18
BiERU (li2022bieru) Bidirectional 64.59
HCAM (dutta2024hcam) Bidirectional 64.4
DialogueGCN (ghosal2019dialoguegcn) Bidirectional 64.18
5 Discussion
Our experiments yield state-of-the-art performance among past-only methods (82.69% for 4-way, 67.07% for 6-way) while revealing several interpretable patterns: hierarchical encoding helps only without context, emotions differ dramatically in context requirements (+8%p to +22%p), and external lexicons provide no benefit. Our discourse marker analysis of 5,286 occurrences uncovers emotion-specific positioning patterns. Below, we discuss four key implications of these findings.
5.1 Conversational Context Subsumes Hierarchical Structure
Our results reveal a striking interaction between encoding strategy and context availability. At utterance-level (K=0), hierarchical encoding outperforms flat encoding by +3.17%p (4-way) and +1.80%p (6-way), confirming that intra-utterance sentence structure provides useful signal for emotion recognition. However, this advantage disappears—and even reverses—when conversational context is incorporated: flat encoding achieves 82.69% versus 81.89% for hierarchical (4-way), and 67.07% versus 66.73% (6-way).
This pattern suggests that turn-level context subsumes the structural information captured by hierarchical encoding. When the model can access preceding utterances, it learns emotional dynamics from the conversation flow, rendering fine-grained sentence boundaries within a single utterance redundant. This finding has practical implications: simpler flat architectures suffice when sufficient conversational context is available, reducing computational overhead without sacrificing performance.
5.2 Do Utterance Peripheries Carry Emotional Information?
Although pooling strategies did not yield statistically significant differences (Friedman $p>.08$), the pattern of best configurations reveals a meaningful connection to discourse structure. At utterance-level (K=0), position-weighted pooling consistently outperformed mean pooling: wmean_pos (emphasizing final tokens) for flat encoding, wmean_pos_rev (emphasizing initial tokens) for hierarchical encoding. With conversational context (K>>0), this advantage disappeared and mean pooling sufficed.
This pattern aligns with our discourse marker analysis. We found a significant association between emotion and peripheral positioning ($\chi^{2}$ test, $p<.0001$, Cramér’s $V=0.062$). Critically, Sad utterances show reduced left-periphery usage (21.9%) compared to Neutral (31.7%), Happy (29.7%), and Angry (28.2%). Post-hoc comparisons confirm Sad differs significantly from all other emotions (all $p<.01$ with Bonferroni correction).
Left-periphery markers like “well” and “oh” signal floor-claiming and turn-management—functions associated with active discourse engagement. The reduced left-periphery usage in Sad utterances suggests that sadness manifests through diminished pragmatic signaling, consistent with the low-arousal, withdrawal-oriented nature of this emotion. When conversational context is unavailable, position-weighted pooling captures these peripheral cues; when context is available, inter-turn dynamics render this positional information redundant.
5.3 Why Does Sadness Require More Context?
Our emotion-specific analysis reveals that Sad benefits most from context (+22.31%p) while Angry benefits least (+8.34%p). An arousal-based explanation would predict that high-arousal emotions are expressed explicitly while low-arousal emotions require contextual inference. However, this account fails: Happy is high-arousal yet benefits substantially from context (+15.82%p), far more than Angry.
We propose that the key factor is not arousal but explicitness of linguistic markers. Angry utterances contain salient lexical cues—profanity, emphatic negation (“I can’t believe”), exclamatory expressions—that are recognizable without conversational history. Sad utterances, in contrast, often lack explicit markers. Expressions like “I see,” “Oh,” or “Yeah, I guess” are pragmatically ambiguous; their emotional valence emerges only from the preceding conversational trajectory.
This interpretation connects to our discourse marker findings: Sad shows reduced use of explicit pragmatic signals (left-periphery markers), forcing the model to rely on accumulated context for disambiguation. Happy presents an intermediate case—positive expressions can be ambiguous with sarcasm or polite neutrality, explaining why context helps despite high arousal. Angry, with its explicit linguistic markers, requires minimal contextual support.
5.4 Is the 6-way Taxonomy Linguistically Justified?
Our confusion matrix analysis raises fundamental questions about the appropriateness of IEMOCAP’s 6-way emotion taxonomy for text-based classification. Contrary to the intuition that emotions of similar valence should be more confusable, we find that Happy–Sad confusion (17.3% turn-level, 18.9% utterance-level) substantially exceeds Happy–Excited confusion (5.3% turn-level, 10.9% utterance-level). This suggests that text-based features do not align with the valence-based groupings assumed by the taxonomy.
The Happy–Excited confusion is also notably asymmetric: 20.6% of Excited samples are misclassified as Happy at utterance-level, while only 1.3% of Happy becomes Excited. This asymmetry suggests that Happy functions as a “catch-all” positive category, absorbing cases that lack the prosodic intensity markers distinguishing Excited. Without acoustic cues, the model defaults to the more frequent positive label.
These findings suggest that the 6-way taxonomy may conflate orthogonal dimensions—valence and arousal—into arbitrary categorical boundaries. For text-only ERC, researchers should consider dimensional models (valence-arousal space) as alternatives, hierarchical classification that separates valence before fine-grained categories, or reduced taxonomies that merge linguistically indistinguishable categories such as Happy and Excited.
6 Limitations and Future Directions
While our findings reveal consistent and interpretable patterns in how models encode emotion, several limitations should be acknowledged and motivate future extensions.
Statistical tendencies rather than deterministic rules.
Although the positional effects we report are statistically reliable, their effect sizes remain modest (Cramér’s $V=0.062$). Emotional expression is inherently variable across speakers and contexts, and our observed patterns should be interpreted as probabilistic tendencies rather than deterministic linguistic laws. Future work should test whether these discourse-level patterns generalize across more diverse corpora such as MELD (poria2018meld), CMU-MOSI (zadeh2016mosi), and CMU-MOSEI (zadeh2018multimodal).
Scope of text-only modeling.
Our analysis is limited to text-based emotion recognition, which captures linguistic and discourse-level cues but omits prosodic and visual signals. Given that emotions like Excited are distinguished primarily through acoustic intensity, extending this framework to multimodal data would enable precise quantification of where and how much information from prosody or facial dynamics contributes beyond text.
Dataset bias and generalizability.
Because IEMOCAP contains acted English dialogues, its label distributions and emotional dynamics may not reflect spontaneous or cross-linguistic behavior. The emotion-specific context patterns we observed (e.g., Sad requiring extended context while Angry saturates quickly) might thus be dataset-specific artifacts. Cross-dataset and multilingual replication will be essential to determine whether these represent general computational principles or culturally contingent effects.
Discourse marker inventory.
Our analysis relies on a predefined 20-marker inventory drawn from established taxonomies. While these markers are well-attested in the linguistics literature, this closed set may miss emotion-specific expressions or informal markers prevalent in conversational speech. Data-driven marker discovery could complement our theory-driven approach.
7 Conclusion
We presented a systematic analysis of emotion recognition in conversation, addressing two gaps in the literature: understanding which architectural choices matter for recognition, and identifying linguistic patterns that could inform generation.
For recognition, our experiments reveal three key findings. First, conversational context is paramount: incorporating preceding dialogue yields substantial improvements, with our past-only approach achieving state-of-the-art results (82.69% for 4-way, 67.07% for 6-way) among text-only methods. Second, hierarchical sentence encoding helps at utterance-level but provides no benefit once conversational context is available, suggesting that turn-level dynamics subsume intra-utterance structure. Third, external affective lexicons (SenticNet) provide no improvement, indicating that pre-trained encoders already capture sufficient emotional semantics.
For linguistic analysis, our examination of 5,286 discourse marker occurrences reveals emotion-specific positioning patterns. Sad utterances show significantly reduced left-periphery marker usage (21.9% vs. 28–32% for other emotions), consistent with diminished turn-management behavior associated with low-arousal withdrawal. This connects to our finding that Sad benefits most from context (+22%p): lacking explicit pragmatic markers, sadness must be inferred from conversational trajectory.
Our confusion analysis further questions the validity of 6-way emotion taxonomies for text-only classification. The asymmetric Happy–Excited confusion (20.6% vs. 1.3%) and unexpectedly high cross-valence errors suggest that categorical boundaries may not align with linguistic distinguishability.
These findings provide actionable insights for ERC system design: simple flat encoders with moderate context (K=30K=305050 preceding turns) capture most predictive information, position-weighted pooling helps only without context, and 4-way classification may be more appropriate for text-only applications. For emotion-conditioned generation, our discourse marker findings suggest concrete production guidelines, such as reducing turn-initial markers for sad utterances.
Appendix A Discourse Marker Inventory
Our discourse marker analysis uses markers drawn from established taxonomies in discourse and pragmatics research. We compiled a search inventory from schiffrin1987discourse, fraser1999discourse, traugott2010intersubjectivity, verhagen2005constructions, aijmer2013understanding, beeching2014discourse, and biber1989styles. Table 6 lists the 20 markers that were empirically found in IEMOCAP, along with their frequencies and theoretical sources.
Table 6: Discourse markers found in IEMOCAP with occurrence counts and source references.
Marker Category Count Source
and Elaborative 2,372 schiffrin1987discourse; fraser1999discourse
so Inferential 1,968 schiffrin1987discourse; fraser1999discourse; beeching2014discourse
like Pragmatic particle 1,210 aijmer2013understanding
but Contrastive 760 schiffrin1987discourse; fraser1999discourse; beeching2014discourse
well Turn-management 727 schiffrin1987discourse; beeching2014discourse
oh Turn-management 564 schiffrin1987discourse; beeching2014discourse
you know Intersubjective 393 schiffrin1987discourse; verhagen2005constructions
i mean Intersubjective 240 schiffrin1987discourse; verhagen2005constructions
maybe Epistemic (doubt) 195 traugott2010intersubjectivity; biber1989styles
though Contrastive 162 fraser1999discourse; beeching2014discourse
i think Epistemic (stance) 131 traugott2010intersubjectivity
probably Epistemic (doubt) 83 traugott2010intersubjectivity; biber1989styles
i guess Epistemic (stance) 77 traugott2010intersubjectivity
yet Contrastive 29 fraser1999discourse
also Elaborative 18 fraser1999discourse
i believe Epistemic (stance) 10 traugott2010intersubjectivity
however Contrastive 6 fraser1999discourse
although Contrastive 5 fraser1999discourse
unfortunately Attitudinal 4 biber1989styles
therefore Inferential 1 fraser1999discourse
Total 8,955
Appendix B Hyperparameter Sensitivity Analysis
To ensure fair comparison, we verified that alternative hyperparameter choices do not yield statistically significant performance differences. We conducted 14 pairwise comparisons across layer selection and pooling methods using paired t-tests and Friedman tests with 10 random seeds.
B.1 Layer Selection: last vs avg_last4
Table 7 presents weighted F1 scores (%) for different layer extraction methods at utterance-level (K=0). None of the six comparisons showed significant differences (all $p>0.24$).
Table 7: Layer comparison at utterance-level (paired t-test, $n=10$)
Task Pooling last avg_last4 $p$-value
4-way mean 64.89 $\pm$ 0.79 64.63 $\pm$ 0.91 0.382
4-way wmean_pos 64.61 $\pm$ 1.06 64.30 $\pm$ 0.94 0.330
4-way wmean_pos_rev 64.65 $\pm$ 1.23 64.54 $\pm$ 0.73 0.753
6-way mean 52.30 $\pm$ 0.91 52.11 $\pm$ 0.81 0.677
6-way wmean_pos 52.42 $\pm$ 1.01 52.05 $\pm$ 1.04 0.508
6-way wmean_pos_rev 52.25 $\pm$ 0.77 51.71 $\pm$ 0.89 0.244
B.2 Pooling Method Comparison
Table 8 shows performance across three pooling methods using the Friedman test. No significant differences were found at either utterance-level or turn-level (all $p>0.08$).
Table 8: Pooling method comparison (Friedman test, $n=10$)
Level Task mean wmean_pos wmean_pos_rev $p$-value
Utterance-level (K=0), layer=last
4-way 64.89 64.61 64.65 0.670
6-way 52.30 52.42 52.25 0.905
Utterance-level (K=0), layer=avg_last4
4-way 64.63 64.30 64.54 0.123
6-way 52.11 52.05 51.71 0.082
Turn-level (best K per seed)
4-way 82.69 $\pm$ 0.50 82.49 $\pm$ 0.46 82.37 $\pm$ 0.67 0.301
6-way 66.88 $\pm$ 0.84 67.07 $\pm$ 0.69 66.57 $\pm$ 0.48 0.150
B.3 Hierarchical Aggregation Comparison
For hierarchical encoding at turn-level, we compared aggregation methods (Table 9). No significant differences were observed (all $p>0.17$).
Table 9: Hierarchical aggregation comparison (paired t-test, $n=10$)
Task mean wmean_pos $t$-statistic $p$-value
4-way 81.89 $\pm$ 0.41 81.57 $\pm$ 0.56 1.39 0.197
6-way 66.73 $\pm$ 0.87 66.19 $\pm$ 0.66 1.47 0.175
B.4 Summary
Table 10 summarizes all 14 comparisons. None showed statistically significant differences ($p<0.05$), justifying our reporting of only the best-performing configurations in the main results.
Table 10: Summary of hyperparameter sensitivity analysis
Category # Tests Significant Min $p$
Layer (last vs avg_last4) 6 0/6 0.244
Pooling (utterance-level) 4 0/4 0.082
Pooling (turn-level FLAT) 2 0/2 0.150
Aggregation (turn-level HIER) 2 0/2 0.175
Total 14 0/14 0.082
Appendix C SenticNet Ablation Studies
Table 11 presents the complete results of SenticNet fusion experiments across 36 configurations.
Table 11: Complete SenticNet Fusion Results (36 Configurations). $\Delta$ shows performance change in percentage points. Each configuration evaluated with 10 seeds.
Encoder Task $\alpha$ Base +Sentic $\Delta$(%) $p$
BERT 4-way 0.05 .633 .628 -0.50 .029
BERT 4-way 0.10 .633 .627 -0.57 .013
BERT 4-way 0.20 .633 .627 -0.58∗∗ .008
BERT 4-way 0.50 .633 .628 -0.49 .022
BERT 4-way 1.00 .633 .628 -0.52∗∗ .007
BERT 4-way concat .633 .628 -0.52∗∗ .007
BERT 6-way 0.05 .485 .482 -0.23 .298
BERT 6-way 0.10 .485 .482 -0.26 .282
BERT 6-way 0.20 .485 .481 -0.35 .146
BERT 6-way 0.50 .485 .483 -0.14 .537
BERT 6-way 1.00 .485 .484 -0.10 .623
BERT 6-way concat .485 .484 -0.10 .623
RoBERTa 4-way 0.05 .634 .633 -0.09 .705
RoBERTa 4-way 0.10 .634 .634 ++0.00 .999
RoBERTa 4-way 0.20 .634 .634 ++0.01 .958
RoBERTa 4-way 0.50 .634 .634 -0.00 .996
RoBERTa 4-way 1.00 .634 .633 -0.07 .804
RoBERTa 4-way concat .634 .633 -0.07 .804
RoBERTa 6-way 0.05 .487 .487 ++0.05 .851
RoBERTa 6-way 0.10 .487 .488 ++0.07 .794
RoBERTa 6-way 0.20 .487 .486 -0.04 .895
RoBERTa 6-way 0.50 .487 .487 ++0.01 .981
RoBERTa 6-way 1.00 .487 .487 ++0.06 .777
RoBERTa 6-way concat .487 .487 ++0.06 .777
S-RoBERTa 4-way 0.05 .656 .653 -0.28 .230
S-RoBERTa 4-way 0.10 .656 .653 -0.32 .231
S-RoBERTa 4-way 0.20 .656 .653 -0.34 .205
S-RoBERTa 4-way 0.50 .656 .653 -0.29 .268
S-RoBERTa 4-way 1.00 .656 .654 -0.25 .385
S-RoBERTa 4-way concat .656 .654 -0.25 .385
S-RoBERTa 6-way 0.05 .516 .516 ++0.08 .794
S-RoBERTa 6-way 0.10 .516 .516 ++0.07 .815
S-RoBERTa 6-way 0.20 .516 .516 ++0.04 .884
S-RoBERTa 6-way 0.50 .516 .515 -0.01 .963
S-RoBERTa 6-way 1.00 .516 .516 ++0.02 .925
S-RoBERTa 6-way concat .516 .516 ++0.02 .925
$^{*}p<.05$, $^{**}p<.01$ (paired $t$-test). Fusion: $\mathbf{e}=(1-\alpha)\mathbf{e}_{\text{ctx}}+\alpha\mathbf{e}_{\text{sentic}}$.