### v2

# Understanding Emotion in Discourse: Recognition Insights and Linguistic Patterns for Generation

Abstract
Despite strong recent progress in Emotion Recognition in Conversation (ERC), two gaps remain: we still lack a clear understanding of which modeling choices materially affect performance, and we have limited linguistic analysis that links recognition findings to actionable cues for generation. We address both gaps via a systematic study on IEMOCAP.
For recognition, we conduct controlled ablations with 10 random seeds and paired tests over seeds (with correction for multiple comparisons), yielding three findings. First, conversational context is the dominant factor: performance saturates quickly, with roughly 90% of the gain observed within our context sweep achieved using only the most recent 10–30 preceding turns (depending on the label set). Second, hierarchical sentence representations improve utterance-only recognition (K=0K{=}0), but the benefit vanishes once turn-level context is available, suggesting that conversational history subsumes much of the intra-utterance structure. Third, a simple integration of an external affective lexicon (SenticNet) does not improve results, consistent with pretrained encoders already capturing much of the affective signal needed for ERC. Under a strictly causal (past-only) setting, our simple models attain strong performance (82.69% 4-way; 67.07% 6-way weighted F1), indicating that competitive accuracy is achievable without access to future turns.
For linguistic analysis, we examine 5,286 discourse-marker occurrences and find a reliable association between emotion and marker position within the utterance ($p<0.0001$). In particular, Sad utterances show reduced left-periphery marker usage (21.9%) relative to other emotions (28–32%), aligning with accounts that link left-periphery markers to active discourse management. This pattern is consistent with our recognition results, where Sad benefits most from conversational context (+22%p), suggesting that sadness often relies more on discourse history than on overt pragmatic signaling in the utterance itself.
1 Introduction
Emotion Recognition in Conversation (ERC) is a key capability for socially intelligent dialogue systems, mental-health support tools, and empathetic conversational agents. Unlike sentence-level emotion classification, ERC requires models to interpret an utterance in relation to conversational history, speaker dynamics, and pragmatic signals that are often subtle or indirect. Recent text-only approaches built on large pretrained encoders have achieved strong performance on benchmarks such as IEMOCAP (Dutta and Ganapathy, 2024), but progress has been accompanied by a growing mismatch between empirical gains and interpretability: it remains unclear which architectural ingredients actually matter and what linguistic evidence these models are exploiting.
Gap 1: Which modeling choices materially matter in modern ERC?
The ERC literature has introduced increasingly elaborate components—hierarchical encoders (Majumder et al., 2019), graph/knowledge-based modules (Zhong et al., 2020; Ghosal et al., 2019), lexicon fusion (Tu et al., 2022), and sophisticated context modeling (Ghosal et al., 2019). However, many reported improvements rely on single-seed runs and heterogeneous experimental setups, making it difficult to separate robust effects from variance or configuration-specific artifacts. As a result, foundational design questions remain unresolved. How much conversational context is truly necessary before gains saturate? Does modeling intra-utterance sentence structure still help once turn-level context is available? Do external affective lexicons add signal beyond what pretrained encoders already capture?
Gap 2: Recognition rarely translates into linguistic guidance for generation.
High recognition accuracy does not automatically reveal how emotions are expressed in discourse, nor does it provide actionable guidance for emotion-conditioned generation. In linguistics, discourse markers are well known to encode stance, subjectivity, and intersubjectivity (Schiffrin, 1987; Beeching and Detges, 2014). Their distribution at the left and right peripheries of utterances is linked to discourse management and listener-oriented effects (Fraser, 1999; Traugott, 2010; Beeching and Detges, 2014). Yet ERC research has rarely examined whether such positional patterns vary systematically by emotion. Bridging recognition and linguistic analysis could benefit both directions: it can provide interpretable hypotheses about what models exploit, and it can yield concrete production cues for generation.
1.1 Research Questions
We address these gaps through a systematic study on IEMOCAP that combines controlled ablation experiments with large-scale discourse-marker analysis. Our investigation is guided by four research questions:
Recognition: What architectural choices matter?
  • RQ1
    Does conversational context improve emotion recognition, and how much context is sufficient before performance saturates?
  • RQ2
    Does hierarchical sentence representation help beyond flat utterance encoding, and under what context conditions?
  • RQ3
    Does incorporating an external affective lexicon (SenticNet) help under a simple integration scheme?
Linguistic Analysis: What patterns exist?
  • RQ4
    Are there emotion-specific discourse-marker positioning patterns that could inform emotion-conditioned generation?
1.2 Contributions
  1. 1.
    We provide a statistically grounded ablation framework for ERC by reporting 10-seed evaluations with paired significance testing and correction for multiple comparisons, enabling reliable component-level conclusions.
  2. 2.
    We isolate the relative contributions of context length, hierarchical structure, and lexicon fusion (Tu et al., 2022), showing that conversational context dominates and saturates quickly, while hierarchical encoding primarily benefits utterance-only settings.
  3. 3.
    We show that emotions differ substantially in their dependence on conversational history, with sad benefiting more from context than angry, suggesting that context effects are not fully explained by arousal alone.
  4. 4.
    We connect recognition to linguistic evidence by analyzing discourse-marker usage and positioning (Schiffrin, 1987; Fraser, 1999; Beeching and Detges, 2014; Traugott, 2010), demonstrating a reliable association between emotion and marker position, including reduced left-periphery usage in sad utterances.
  5. 5.
    Under strictly causal (past-only) constraints, our simple models achieve strong performance on both 4-way and 6-way classification, illustrating that competitive accuracy is attainable without access to future turns—a practical consideration for real-time deployment.
2 Related Work
2.1 Emotion Recognition in Conversation
Early ERC methods focused on capturing conversational dynamics through recurrent architectures. DialogueRNN (Majumder et al., 2019) models speaker states across turns, achieving 76.2% on IEMOCAP 4-way classification. COSMIC (Ghosal et al., 2020) incorporates commonsense knowledge for context enhancement (77.4%). Recent work has pursued increasingly complex architectures, including graph-based models (Ghosal et al., 2019), hierarchical attention (Ma et al., 2022), and multimodal transformers (Hu et al., 2022).
A central axis of variation across ERC systems is how conversational context is operationalized. Several approaches adopt heuristic context windows by concatenating neighboring utterances as raw text; for example, UniMSE concatenates the current utterance with a symmetric window of two preceding and two following turns, which implicitly relies on future context (Hu et al., 2022). Other models assume a fixed number of preceding turns as the local interaction neighborhood (e.g., t=3t{=}3) (Mai et al., 2019), treating context length as a default hyperparameter rather than an empirically justified operating point. For transformer baselines, long contexts are often handled by concatenation with truncation: when the input exceeds the model’s length limit, remote utterances are discarded, imposing an implicit cutoff on usable history (e.g., Shen et al. (2021)).
Among high-performing text-only methods, HCAM (Dutta and Ganapathy, 2024) explicitly introduces inter-utterance context via a bidirectional GRU (Bi-GRU) with self-attention, aiming to leverage information from the full conversation. Their formulation uses Bi-GRU to incorporate contextual signals and adds self-attention to better handle long-range dependencies across dialogues with many utterances. While effective, such bidirectional context access is not available in real-time settings where future turns are unknown.
In contrast, we focus on strictly causal (past-only) modeling and explicitly characterize the performance–context trade-off by sweeping the number of preceding turns and quantifying saturation behavior. This complements prior work by replacing heuristic context choices (fixed windows or implicit truncation) with a systematic analysis of how much past text is sufficient, and whether context requirements differ across emotions.
2.2 Discourse Markers and Affective Lexicons
Linguistic theory has long recognized discourse markers as signals of speaker stance and emotion (Schiffrin, 1987). Fraser (1999) classified pragmatic markers by function, while Beeching and Detges (2014) demonstrated their tendency to cluster at utterance peripheries. The left periphery often hosts markers expressing speaker stance or discourse management (e.g., well signaling hesitation, oh marking surprise or realization, and actually introducing contrast or correction), while right-peripheral markers can be more hearer-oriented (Traugott, 2010; Beeching and Detges, 2014).
Despite extensive theoretical study, computational ERC models have rarely examined whether such positional patterns vary systematically by emotion. We provide a large-scale corpus analysis of discourse-marker positioning in IEMOCAP and evaluate its association with emotion labels, complementing prior linguistic accounts of peripheral distributions.
Affective lexicons provide complementary emotional knowledge. While early resources such as WordNet-Affect (Strapparava, 2004) offered categorical labels, SenticNet (Cambria et al., 2024) maps concepts onto psychological dimensions. Knowledge-enhanced models (Zhong et al., 2020; Tu and others, 2022) report gains, but whether lexicons add unique information beyond contextual encoders remains unclear—a question we examine empirically under a simple integration scheme.
3 Methodology
(a) Dialogue length distribution
(b) Sentence distribution
(c) 4-way emotion patterns
(d) 6-way emotion patterns
Figure 1: IEMOCAP dataset characteristics.
3.1 Task and Dataset
We conduct experiments on the IEMOCAP dataset (Busso et al., 2008), which contains approximately 12 hours of audiovisual data from 10 actors performing scripted and improvised dyadic conversations across 110 dialogue sessions. The dataset exhibits substantial variability in dialogue length (Figure 1a), with sessions ranging from 24 to 267 utterances (mean: 91.7, median: 67.5). Approximately 70% of dialogues contain fewer than 100 utterances, while the 95th percentile reaches 196 utterances, highlighting the long-tail distribution inherent in conversational data.
Turn-level context length and $K$-sweep protocol.
We define the context length $K$ as the number of preceding turns included as strictly past-only conversational history for a target utterance. Because there is no established stopping rule for how much prior dialogue is sufficient in ERC, we perform a bottom-up sweep $K\in\{0,1,\ldots,K_{\max}\}$ to characterize the performance–context trade-off and identify saturation behavior. We set $K_{\max}=200$ based on the dialogue-length distribution (covering the long tail while keeping the sweep computationally tractable). [^1] The maximum dialogue length in IEMOCAP is 267 turns; our choice of $K_{\max}=200$ is driven by the empirical distribution (Figure 1a), where 200 exceeds the 95th percentile, and is used for analysis rather than model selection. Crucially, the sweep is used strictly for analysis (saturation characterization) and not for test-set model selection; we do not choose $K$ to maximize test performance.
When a target utterance has fewer than $K$ preceding turns (e.g., early in a dialogue), we use all available history (a variable-length prefix). For turn-level models, sequences are left-padded for batching and padded positions are masked so they do not contribute to contextual aggregation.
For a fair comparison with state-of-the-art models (Dutta and Ganapathy, 2024), we follow the standard speaker-disjoint split strategy: Sessions 2–4 serve as training data (6,072 utterances from 68 dialogues), Session 1 as validation (1,819 utterances from 20 dialogues), and Session 5 as test set (2,196 utterances from 22 dialogues). This split ensures speaker independence between training and evaluation, with each dialogue containing an average of 89–100 utterances across splits. The sentence-level structure within utterances (Figure 1b) shows that most utterances contain 2–5 sentences, motivating our comparison of flat versus hierarchical encoding strategies. Sentence segmentation for hierarchical encoding is performed with spaCy’s sentencizer after lowercasing and standard token normalization. [^2] We follow common ERC preprocessing practice (lowercasing and tokenization) and use a deterministic sentence boundary detector to ensure reproducibility.
We evaluate on two emotion taxonomies commonly used in the literature. The 4-way classification task considers angry (1,103 utterances, 19.9%), happy (1,636, 29.6%), sad (1,084, 19.6%), and neutral (1,708, 30.9%) emotions, where the excited category is merged into happy (Figure 1c). The 6-way task extends this to include excited (1,041, 14.1%) and frustrated (1,849, 25.1%) as separate categories (Figure 1d). We use weighted F1-score as our primary metric to account for class imbalance inherent in conversational emotion data, and we also report class-wise F1 scores and confusion matrices for error analysis.
3.2 Discourse Marker Analysis
Beyond recognition accuracy, we conduct a linguistic analysis of discourse markers (DMs) to examine emotion-specific pragmatic patterns that may inform future work on emotion-conditioned dialogue generation. DMs are lexical expressions that signal relationships between discourse segments rather than contributing to propositional content (Schiffrin, 1987; Fraser, 1999). Drawing from established taxonomies (Schiffrin, 1987; Fraser, 1999; Traugott, 2010; Beeching and Detges, 2014), we identify 20 markers occurring in IEMOCAP, including turn-management markers (well, oh), connectives (and, but, so), and stance markers (I think, I guess, maybe, you know, I mean). See Appendix A for the full inventory and frequency distribution.
For each marker occurrence, we record its relative position within the utterance (normalized to $[0,1]$), its periphery classification, and the emotion label of the containing utterance. Following standard discourse accounts of utterance peripheries (Beeching and Detges, 2014), we operationally define the left periphery (LP) as position <0.15<0.15, the right periphery (RP) as position >0.85>0.85, and medial otherwise. For consistency with our pooling notation, wmean_pos_rev emphasizes utterance-initial (left-peripheral) tokens, whereas wmean_pos emphasizes utterance-final (right-peripheral) tokens; mean pooling treats all positions uniformly.
To test for emotion-specific positional patterns, we employ (i) ANOVA to compare mean positions across emotions, (ii) $\chi^{2}$ tests with Cramér’s $V$ to assess association between periphery categories and emotions, and (iii) mixed-effects models with dialogue as a random intercept to control for dialogue-level variation (e.g., $\text{pos}\sim\text{emotion}+(1|\text{dialogue})$). We apply post-hoc pairwise comparisons with Bonferroni correction where appropriate.
[…]
Figure 2: Model architecture. Each turn is encoded independently via Sentence-RoBERTa using either flat (whole utterance) or hierarchical (sentence-level) encoding. SenticNet features are optionally fused (dashed boxes). For classification, we use MLP when K=0 (no context) or unidirectional LSTM when K>0 (with preceding turns as context).
3.3 Model Architecture
To investigate how emotional information is encoded and propagated in dialogue, we develop two encoder variants: a flat encoder and a hierarchical encoder. The flat encoder processes each utterance as a single sequence, while the hierarchical encoder first encodes individual sentences within an utterance, then aggregates them to form the utterance representation. Figure 2 illustrates our architecture.
Utterance Encoder.
We compare (1) flat encoding, which treats each utterance as a single sequence and extracts a pooled representation, and (2) hierarchical encoding, which encodes each sentence and then aggregates sentence representations into an utterance vector. We evaluate three pre-trained encoders: BERT-base-uncased (Devlin et al., 2019), RoBERTa-base (Liu et al., 2019), and Sentence-RoBERTa (NLI-RoBERTa-base-v2) (Reimers and Gurevych, 2019). Unless stated otherwise, encoders are used as fixed feature extractors and utterance embeddings are precomputed to isolate the effects of contextual modeling and pooling choices. [^3] This design avoids confounding the analysis with end-to-end fine-tuning instability across seeds.
We compare two layer strategies: (i) avg_last4, averaging the last four transformer layers, and (ii) last, using only the final layer output. We use a maximum input length of 128 subword tokens for the utterance encoder; longer utterances are truncated from the left/right as in standard transformer preprocessing.
Context construction and label availability.
During embedding generation, we include all turns in each dialogue in chronological order, including utterances without target emotion labels or outside the target label set. These turns may appear in the contextual history, but only target-labeled utterances contribute to the training loss and evaluation metrics. This allows long-range history to be represented without altering the supervised label space.
Classifier.
For utterance-level classification (K=0K{=}0), the utterance representation is passed through a two-layer MLP with ReLU activation and dropout. When incorporating turn-level context (K>0K{>}0), we process the sequence of utterance representations via a single-layer unidirectional LSTM and use the final hidden state for classification.
Lexical Feature Integration.
To examine whether external affective knowledge benefits emotion recognition, we integrate SenticNet 7 (Cambria et al., 2024), which provides four-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude) for words and phrases. We lowercase and tokenize each utterance, match SenticNet entries at the token level (and, where available, multiword expressions), and aggregate matched vectors via mean pooling; unmatched tokens contribute zeros. The resulting 4D vector is concatenated with the encoder representation.
3.4 Training and Implementation Details
Our classifier consists of a two-layer MLP (for K=0K{=}0) or a single-layer unidirectional LSTM (for K>0K{>}0). We use fixed hyperparameters across all experiments: learning rate $1\mathrm{e}{-3}$, hidden dimension 256, dropout 0.3, and batch size 64. Hyperparameters were selected once on the validation set and then held fixed across all ablations to ensure comparability.
Training employs Adam with early stopping on validation weighted F1 (patience 60 for utterance-level and 20 for turn-level experiments). [^4] We use validation weighted F1 to align the early-stopping criterion with the primary evaluation metric. All experiments use 10 random seeds $\{42,43,\ldots,51\}$ with results reported as mean $\pm$ standard deviation. Statistical comparisons across configurations use paired $t$-tests over seeds with Bonferroni correction where applicable. Implementation uses PyTorch 1.13, and experiments were conducted on NVIDIA A100 GPUs via Saturn Cloud.
3.5 Encoder Selection
We compared three pre-trained text encoders at utterance level (K=0K{=}0) on IEMOCAP 4-way classification (Table 1). Sentence-RoBERTa yields the highest mean weighted F1 (65.29%), which is plausibly attributable to its NLI fine-tuning that produces robust sentence-level semantics. We emphasize, however, that this ranking may be dataset-dependent: IEMOCAP consists of relatively well-structured, acted dyadic dialogues, and noisier, multi-party corpora (e.g., MELD) may favor different encoders. To isolate architectural effects under a fixed encoder, all subsequent experiments use Sentence-RoBERTa unless otherwise noted.
Table 1: Encoder comparison on 4-way classification (K=0, 10 seeds).
Encoder WF1 (%) Std Min Max 95% CI
BERT-base 63.99 0.85 62.62 65.18 [63.38, 64.59]
RoBERTa-base 65.01 0.80 63.90 66.31 [64.44, 65.58]
Sentence-RoBERTa 65.29 1.17 64.10 67.31 [64.45, 66.12]
3.6 Main Results
We report two complementary views of performance. First, for a fair comparison across design choices, we evaluate each configuration at fixed context sizes $K$ (Table 2), where $K$ denotes the number of strictly preceding turns (past-only context). This controlled evaluation shows that conversational context dominates performance: most of the total gain is already realized with a short history (e.g., K=10K{=}10), and performance changes only marginally beyond K=30K{=}305050 depending on the task.
Within the same $K$, neither hierarchical encoding nor position-weighted pooling shows a consistent advantage. In particular, FLAT vs. HIER differences are not statistically significant in the fixed-$K$ setting (paired tests across seeds; p>0.9p{>}0.9), and pooling choices do not yield reliable improvements (all comparisons p>0.08p{>}0.08). These results suggest that once modest conversational history is provided, architectural sophistication at the utterance level contributes little relative to context.
Second, to characterize the performance–context trade-off and identify saturation behavior, we also conduct a full $K$-sweep analysis (Section 3.7). For completeness, we additionally report the best-performing configurations observed in the sweep (Table LABEL:tab:main-results); importantly, these are used for analysis rather than test-set model selection.
[…]
Task Type Pooling K=0K{=}0 K=10K{=}10 K=30K{=}30 K=50K{=}50
[…]
Task Type Pooling K=0K{=}0 K=10K{=}10 K=30K{=}30 K=50K{=}50
[…]
Task Type K=100K{=}100 K=200K{=}200
[…]
6-way HIER (wmean_pos_rev) 62.49 63.74
3.7 Emotion-Specific Context Effects
While prior work on variable-length context focuses on how to adaptively select context windows through speaker-aware modules (Zhang et al., 2023), we take a complementary approach: we systematically investigate what patterns emerge when varying context length and whether different emotions exhibit distinct context requirements.
To analyze emotion-specific effects, we compute per-class F1 scores at each context length from K=0K{=}0 (utterance only) to K=200K{=}200 (preceding turns). For each emotion, we report the context length at which F1 attains its maximum within the sweep range, and quantify improvement relative to K=0K{=}0. We also determine the saturation point, defined as the minimum $K$ at which 90% of the maximum improvement (within the sweep range) is achieved. All analyses are conducted across 10 random seeds.
Our first finding is that performance saturates rapidly in aggregate: a short history (e.g., K=10K{=}10) already recovers most of the eventual gain in weighted F1, and improvements beyond K=30K{=}305050 are modest for the average utterance (Table 2). This suggests that immediate preceding context carries the majority of predictive information for ERC under strictly causal access.
Our second finding reveals a dissociation between saturation timing and improvement magnitude at the emotion level. While emotions reach saturation at broadly similar context lengths (Kruskal–Wallis $H=0.52$, $p=0.91$), the magnitude of improvement differs substantially (one-way ANOVA $F=136.80$, $p<0.0001$). Sad benefits the most from context (+22%p), whereas Angry benefits the least (+8–9%p), and this ordering remains consistent across 4-way and 6-way settings.
Figure 3: Per-emotion F1 scores across context sizes (4-way classification). Angry peaks with relatively short context ($K^{*}\approx 10$), while Sad/Happy/Neutral peak at substantially larger $K$ values. Shaded regions indicate $\pm 1$ standard deviation across 10 seeds.
Finally, although aggregate performance plateaus quickly, the context length at which a given emotion attains its peak within the sweep range can be much larger. This long-tail behavior motivates reporting both fixed-$K$ comparisons (for fairness) and sweep-based analyses (for saturation characterization). We discuss theoretical implications in Section 4.
3.8 Ablation Studies
We conducted ablations over (i) pooling strategy, (ii) external lexical knowledge, and (iii) layer aggregation, using the same training protocol and 10-seed evaluation as in the main experiments.
Pooling strategy.
We compared mean pooling, position-weighted pooling emphasizing utterance-final tokens (wmean_pos), and position-weighted pooling emphasizing utterance-initial tokens (wmean_pos_rev). Across seeds, no pooling method was consistently superior (Friedman test, $p>.08$). In terms of best-performing configurations, position-weighted pooling more often appeared among the top runs at utterance-level (K=0K{=}0), whereas mean pooling was more frequently competitive when conversational context was included (K>0K{>}0); however, these differences were not statistically reliable.
External lexical knowledge.
We augmented utterance representations with SenticNet 7 (Cambria et al., 2024), which provides four affective dimensions (pleasantness, attention, sensitivity, aptitude). This augmentation did not improve performance and in some settings slightly reduced it (e.g., 4-way: $\approx-0.94$ F1 points; 6-way: near-zero change), suggesting that the pretrained encoder representations already capture the affective information that SenticNet provides. Full results across all configurations are reported in Appendix C.
Layer aggregation.
We compared averaging the last four transformer layers (avg_last4) against using only the final layer (last). We observed no significant differences (paired $t$-tests; minimum $p=.244$ across comparisons).
3.9 Comparison with Prior Work
We situate our results against prior text-only ERC methods on IEMOCAP. A key axis of comparison is temporal access: bidirectional models can use both past and future utterances, whereas past-only (causal) models are restricted to preceding context and are therefore deployable in real-time settings. Importantly, prior results are reported under heterogeneous protocols (e.g., context windows, architectures, and often single-run reporting), so the tables below provide a reference comparison rather than a strict head-to-head evaluation.
4-way classification.
Table 3 shows that our past-only model is competitive with, and in some cases exceeds, reported performance of representative text-only baselines, including several bidirectional systems. We report the mean across 10 seeds for statistical robustness, whereas most prior work reports a single number.
[…]
Method Context WF1 (%)
Ours (mean over 10 seeds) Past-only 82.69
HFFN (Mai et al., 2019) Bidirectional 81.54
HCAM (Dutta and Ganapathy, 2024) Bidirectional 81.4
CHFusion (Majumder et al., 2018) Bidirectional 73.6
6-way classification.
On the 6-way task (Table 4), our past-only model achieves 67.07% weighted F1 (mean over 10 seeds), which is comparable to strong reported baselines and exceeds several bidirectional systems. We stress that differences should be interpreted cautiously due to protocol mismatch and the prevalence of single-run reporting in prior work.
Table 4: IEMOCAP 6-way (text-only, excluding LLM-based approaches). Our result is mean over 10 seeds; prior work numbers are reported. Bold indicates best reported past-only result among non-LLM baselines.
Method Context WF1 (%)
EmoCaps (Li et al., 2022c) Bidirectional 69.49
DAG-ERC (Shen et al., 2021) Past-only 68.03
Ours (mean over 10 seeds) Past-only 67.07
SKAIG (Li et al., 2021) Bidirectional 66.96
DialogueCRN (Hu et al., 2021) Bidirectional 66.20
CoG-BART (Li et al., 2022a) Bidirectional 66.18
BiERU (Li et al., 2022b) Bidirectional 64.59
HCAM (Dutta and Ganapathy, 2024) Bidirectional 64.4
DialogueGCN (Ghosal et al., 2019) Bidirectional 64.18
4 Discussion
Our study delivers strong performance under a strictly causal (past-only) setting (82.69% for 4-way; 67.07% for 6-way) and yields several interpretable patterns. Across analyses, conversational context provides the dominant gain, emotion classes differ markedly in their reliance on context, and external lexicons (SenticNet) do not improve performance. In addition, our corpus-based analysis of 5,286 discourse-marker (DM) occurrences reveals small but reliable emotion-specific positional tendencies. Below, we discuss four implications.
4.1 Conversational Context Reduces the Marginal Utility of Hierarchical Structure
We observe an interaction between encoding strategy and context availability. In the best-performing configurations at K=0K{=}0, hierarchical encoding exceeds flat encoding (Table LABEL:tab:main-results), consistent with the intuition that intra-utterance sentence structure can provide additional cues when the model cannot consult surrounding turns.
However, once past conversational context is incorporated, this advantage diminishes: flat and hierarchical variants become closely matched, and the best-performing contextual models are not reliably separated by the choice of utterance encoder. This suggests that turn-level context can partly substitute for fine-grained intra-utterance structure, because emotional evidence is often distributed across adjacent turns (e.g., responses, clarifications, escalation/de-escalation). Practically, this supports using simpler flat encoders in causal, real-time deployments when sufficient past context is available.
4.2 Do Utterance Peripheries Carry Emotion-Relevant Signal?
Pooling variants did not yield statistically significant differences in our multi-seed comparisons (Friedman $p>.08$), yet we consistently observed that position-weighted pooling is competitive at K=0K{=}0, while simple mean pooling is sufficient once context is introduced. We interpret this as a weak but suggestive pattern rather than a definitive effect.
This interpretation is compatible with our discourse-marker analysis. We find a significant association between emotion and DM periphery category ($\chi^{2}$, $p<.0001$), although the effect size is small (Cramér’s $V=0.062$). Notably, Sad utterances show reduced left-periphery usage (21.9%) compared to Neutral (31.7%), Happy (29.7%), and Angry (28.2%), with post-hoc tests indicating that Sad differs from the other emotions (Bonferroni-corrected, all $p<.01$). One plausible account is that left-periphery markers (e.g., well, oh) often participate in turn-management and stance negotiation; their reduced use may reflect more muted pragmatic signaling in Sad speech. Importantly, we treat this as a linguistic hypothesis supported by corpus statistics, not as a proven mechanism of the pooling variants.
4.3 Why Does Sadness Benefit Most from Context?
Emotion-specific analysis shows that Sad gains the most from added context (+22.31%p), whereas Angry gains the least (+8.34%p). A pure arousal-based account is insufficient, because Happy (often high-arousal) also benefits substantially from context (+15.82%p). Instead, our results suggest that what matters is the availability of explicit lexical/pragmatic cues in the target utterance.
Angry turns often contain salient lexical signals (e.g., emphatic negation, direct accusations, profanity) that are informative even without prior history. By contrast, Sad turns can be lexically understated and pragmatically ambiguous (e.g., I see, yeah, I guess), making their emotional interpretation contingent on the preceding trajectory. This aligns with the DM finding that Sad shows reduced left-periphery marking, potentially reducing overt discourse-management signals and increasing reliance on conversational history for disambiguation. More broadly, this supports the view that causal ERC should be evaluated not only by overall F1 but also by how different emotions depend on context.
4.4 Is the 6-way Taxonomy Well-Identified in Text-Only ERC?
Our confusion patterns raise questions about how well the 6-way categories are separable from text alone. In particular, Happy–Sad confusion is higher than Happy–Excited confusion in our analysis, and Happy absorbs a substantial portion of Excited instances, yielding an asymmetric confusion pattern. This is consistent with the interpretation that acoustic intensity cues—often crucial for distinguishing Excited from Happy—are absent in text-only settings, so the model defaults to the more frequent positive label.
These observations do not imply that the 6-way taxonomy is intrinsically invalid; rather, they suggest that in text-only ERC, some category boundaries may be weakly identified. This motivates alternative formulations, such as (i) hierarchical classification (e.g., valence first, then finer labels), (ii) dimensional prediction (valence/arousal), or (iii) reduced label sets that merge categories that are difficult to separate without prosody.
5 Limitations and Future Directions
While our analyses reveal consistent and interpretable patterns in how causal ERC models leverage context and discourse cues, several limitations should be acknowledged and motivate future extensions.
Statistical tendencies rather than deterministic rules.
The discourse-marker effects we report are statistically reliable but small in magnitude (e.g., Cramér’s $V=0.062$). Emotional expression is highly variable across speakers, interactional goals, and situations; accordingly, our positional patterns should be interpreted as probabilistic tendencies rather than deterministic linguistic laws. Future work should quantify robustness under domain shift and across interaction types (e.g., multi-party chat in MELD (Poria et al., 2018), written dialogue in DailyDialog, and other conversational corpora), and should report effect sizes alongside significance to avoid overinterpretation.
Scope of text-only modeling.
Our experiments are restricted to text-based ERC. This captures lexical and discourse-level cues but omits prosodic and visual signals that are often crucial for separating closely related categories (e.g., excited vs. happy). Extending the same ablation and fixed-context protocol to multimodal encoders would allow precise attribution of what additional information is contributed by acoustics and facial dynamics beyond text.
Dataset bias, label noise, and generalizability.
IEMOCAP consists of acted English dyadic dialogues, which may differ from spontaneous interaction in turn-taking style, emotional intensity, and marker usage. In addition, categorical emotion labels in conversational corpora are inherently noisy and sometimes underspecified for text-only interpretation. The emotion-specific context patterns we observe (e.g., larger gains for Sad than Angry) may therefore partially reflect dataset- and annotation-specific properties. Cross-dataset replication and multilingual evaluation are essential to determine whether these are general computational regularities or corpus-contingent effects.
Operationalization choices in context and discourse analysis.
Our context analysis uses turn-based windows and a fixed definition of “past-only” context, and our discourse analysis relies on normalized token positions with predefined periphery thresholds (LP <0.15<0.15, RP >0.85>0.85). Alternative operationalizations (e.g., sentence- or time-based windows, speaker-conditioned windows, syntactic peripheries, or discourse-unit segmentation) may yield different quantitative estimates. Future work should stress-test these choices and assess whether conclusions hold under reasonable variants.
Closed discourse-marker inventory.
We analyze a predefined inventory of 20 markers drawn from established taxonomies. While theory-driven, this closed set may miss informal and multiword pragmatic expressions prevalent in conversational speech. A promising direction is to complement this with data-driven discovery (e.g., collocation-based mining, pragmatic phrase induction, or supervised taggers) and to examine whether automatically discovered marker families show stronger or more generalizable emotion associations.
6 Conclusion
We presented a systematic analysis of emotion recognition in conversation (ERC) that targets two gaps in the literature: (i) clarifying which modeling choices materially affect recognition under a strictly causal (past-only) setting, and (ii) identifying interpretable discourse patterns that can inform future work on emotion-conditioned generation.
For recognition, our experiments yield three primary takeaways. First, conversational context is the dominant driver of performance: most of the attainable gain is achieved with a relatively short window of preceding turns, after which returns diminish. Second, intra-utterance hierarchical sentence encoding can be beneficial when no conversational context is available, but its advantage largely disappears once turn-level context is introduced, suggesting that inter-turn dynamics can substitute for fine-grained within-utterance structure in causal ERC. Third, integrating an external affective lexicon (SenticNet) provides no improvement in our setting, consistent with the view that modern pretrained encoders already capture much of the affective semantics needed for classification.
For linguistic analysis, our corpus study of 5,286 discourse-marker occurrences reveals small but reliable emotion-specific positional tendencies. In particular, Sad utterances exhibit reduced left-periphery marker usage relative to other emotions, which we interpret as diminished overt turn-management signaling. This observation complements the recognition-side finding that Sad benefits most from added context: when explicit pragmatic cues are weaker, conversational history becomes more important for disambiguation.
Finally, our confusion analysis highlights that, in text-only ERC, some fine-grained distinctions in the 6-way taxonomy are weakly identified without prosodic information (e.g., asymmetric Happy–Excited confusions). Taken together, these results suggest practical guidance for causal ERC system design: simple flat encoders with moderate past context can capture most predictive information; more elaborate utterance structure offers limited marginal benefit once context is available; and careful attention to discourse phenomena offers a principled path toward interpretability and generation-relevant insights.
Acknowledgements
This work was supported by the UC Noyce Initiative. We thank the anonymous reviewers for their helpful comments. Portions of the writing and visualizations were developed with assistance from Claude (Anthropic). The authors remain fully responsible for all content.
Appendix A Discourse Marker Inventory
Our discourse marker analysis uses markers drawn from established taxonomies in discourse and pragmatics research. We compiled a search inventory from Schiffrin (1987), Fraser (1999), Traugott (2010), Verhagen (2005), Aijmer (2013), Beeching and Detges (2014), and Biber and Finegan (1989). Table 5 lists the 20 markers that were empirically found in IEMOCAP, along with their frequencies and theoretical sources.
Table 5: Discourse markers found in IEMOCAP with occurrence counts and source references.
Marker Category Count Source
and Elaborative 2,372 Schiffrin (1987); Fraser (1999)
so Inferential 1,968 Schiffrin (1987); Fraser (1999); Beeching and Detges (2014)
like Pragmatic particle 1,210 Aijmer (2013)
but Contrastive 760 Schiffrin (1987); Fraser (1999); Beeching and Detges (2014)
well Turn-management 727 Schiffrin (1987); Beeching and Detges (2014)
oh Turn-management 564 Schiffrin (1987); Beeching and Detges (2014)
you know Intersubjective 393 Schiffrin (1987); Verhagen (2005)
i mean Intersubjective 240 Schiffrin (1987); Verhagen (2005)
maybe Epistemic (doubt) 195 Traugott (2010); Biber and Finegan (1989)
though Contrastive 162 Fraser (1999); Beeching and Detges (2014)
i think Epistemic (stance) 131 Traugott (2010)
probably Epistemic (doubt) 83 Traugott (2010); Biber and Finegan (1989)
i guess Epistemic (stance) 77 Traugott (2010)
yet Contrastive 29 Fraser (1999)
also Elaborative 18 Fraser (1999)
i believe Epistemic (stance) 10 Traugott (2010)
however Contrastive 6 Fraser (1999)
although Contrastive 5 Fraser (1999)
unfortunately Attitudinal 4 Biber and Finegan (1989)
therefore Inferential 1 Fraser (1999)
Total 8,955
Appendix B Hyperparameter Sensitivity Analysis
To ensure fair comparison, we verified that alternative hyperparameter choices do not yield statistically significant performance differences. We conducted 14 pairwise comparisons across layer selection and pooling methods using paired t-tests and Friedman tests with 10 random seeds.
B.1 Layer Selection: last vs avg_last4
Table 6 presents weighted F1 scores (%) for different layer extraction methods at utterance-level (K=0). None of the six comparisons showed significant differences (all $p>0.24$).
Table 6: Layer comparison at utterance-level (paired t-test, $n=10$)
Task Pooling last avg_last4 $p$-value
4-way mean 64.89 $\pm$ 0.79 64.63 $\pm$ 0.91 0.382
4-way wmean_pos 64.61 $\pm$ 1.06 64.30 $\pm$ 0.94 0.330
4-way wmean_pos_rev 64.65 $\pm$ 1.23 64.54 $\pm$ 0.73 0.753
6-way mean 52.30 $\pm$ 0.91 52.11 $\pm$ 0.81 0.677
6-way wmean_pos 52.42 $\pm$ 1.01 52.05 $\pm$ 1.04 0.508
6-way wmean_pos_rev 52.25 $\pm$ 0.77 51.71 $\pm$ 0.89 0.244
B.2 Pooling Method Comparison
Table 7 shows performance across three pooling methods using the Friedman test. No significant differences were found at either utterance-level or turn-level (all $p>0.08$).
Table 7: Pooling method comparison (Friedman test, $n=10$)
Level Task mean wmean_pos wmean_pos_rev $p$-value
Utterance-level (K=0), layer=last
4-way 64.89 64.61 64.65 0.670
6-way 52.30 52.42 52.25 0.905
Utterance-level (K=0), layer=avg_last4
4-way 64.63 64.30 64.54 0.123
6-way 52.11 52.05 51.71 0.082
Turn-level (best K per seed)
4-way 82.69 $\pm$ 0.50 82.49 $\pm$ 0.46 82.37 $\pm$ 0.67 0.301
6-way 66.88 $\pm$ 0.84 67.07 $\pm$ 0.69 66.57 $\pm$ 0.48 0.150
B.3 Hierarchical Aggregation Comparison
For hierarchical encoding at turn-level, we compared aggregation methods (Table 8). No significant differences were observed (all $p>0.17$).
Table 8: Hierarchical aggregation comparison (paired t-test, $n=10$)
Task mean wmean_pos $t$-statistic $p$-value
4-way 81.89 $\pm$ 0.41 81.57 $\pm$ 0.56 1.39 0.197
6-way 66.73 $\pm$ 0.87 66.19 $\pm$ 0.66 1.47 0.175
B.4 Summary
Table 9 summarizes all 14 comparisons. None showed statistically significant differences ($p<0.05$), justifying our reporting of only the best-performing configurations in the main results.
Table 9: Summary of hyperparameter sensitivity analysis
Category # Tests Significant Min $p$
Layer (last vs avg_last4) 6 0/6 0.244
Pooling (utterance-level) 4 0/4 0.082
Pooling (turn-level FLAT) 2 0/2 0.150
Aggregation (turn-level HIER) 2 0/2 0.175
Total 14 0/14 0.082
Appendix C SenticNet Ablation Studies
Table 10 presents the complete results of SenticNet fusion experiments across 36 configurations.
Table 10: Complete SenticNet Fusion Results (36 Configurations). $\Delta$ shows performance change in percentage points. Each configuration evaluated with 10 seeds.
Encoder Task $\alpha$ Base +Sentic $\Delta$(%) $p$
BERT 4-way 0.05 .633 .628 -0.50 .029
BERT 4-way 0.10 .633 .627 -0.57 .013
BERT 4-way 0.20 .633 .627 -0.58∗∗ .008
BERT 4-way 0.50 .633 .628 -0.49 .022
BERT 4-way 1.00 .633 .628 -0.52∗∗ .007
BERT 4-way concat .633 .628 -0.52∗∗ .007
BERT 6-way 0.05 .485 .482 -0.23 .298
BERT 6-way 0.10 .485 .482 -0.26 .282
BERT 6-way 0.20 .485 .481 -0.35 .146
BERT 6-way 0.50 .485 .483 -0.14 .537
BERT 6-way 1.00 .485 .484 -0.10 .623
BERT 6-way concat .485 .484 -0.10 .623
RoBERTa 4-way 0.05 .634 .633 -0.09 .705
RoBERTa 4-way 0.10 .634 .634 ++0.00 .999
RoBERTa 4-way 0.20 .634 .634 ++0.01 .958
RoBERTa 4-way 0.50 .634 .634 -0.00 .996
RoBERTa 4-way 1.00 .634 .633 -0.07 .804
RoBERTa 4-way concat .634 .633 -0.07 .804
RoBERTa 6-way 0.05 .487 .487 ++0.05 .851
RoBERTa 6-way 0.10 .487 .488 ++0.07 .794
RoBERTa 6-way 0.20 .487 .486 -0.04 .895
RoBERTa 6-way 0.50 .487 .487 ++0.01 .981
RoBERTa 6-way 1.00 .487 .487 ++0.06 .777
RoBERTa 6-way concat .487 .487 ++0.06 .777
S-RoBERTa 4-way 0.05 .656 .653 -0.28 .230
S-RoBERTa 4-way 0.10 .656 .653 -0.32 .231
S-RoBERTa 4-way 0.20 .656 .653 -0.34 .205
S-RoBERTa 4-way 0.50 .656 .653 -0.29 .268
S-RoBERTa 4-way 1.00 .656 .654 -0.25 .385
S-RoBERTa 4-way concat .656 .654 -0.25 .385
S-RoBERTa 6-way 0.05 .516 .516 ++0.08 .794
S-RoBERTa 6-way 0.10 .516 .516 ++0.07 .815
S-RoBERTa 6-way 0.20 .516 .516 ++0.04 .884
S-RoBERTa 6-way 0.50 .516 .515 -0.01 .963
S-RoBERTa 6-way 1.00 .516 .516 ++0.02 .925
S-RoBERTa 6-way concat .516 .516 ++0.02 .925
$^{*}p<.05$, $^{**}p<.01$ (paired $t$-test). Fusion: $\mathbf{e}=(1-\alpha)\mathbf{e}_{\text{ctx}}+\alpha\mathbf{e}_{\text{sentic}}$.